# Learning to Generate Instruction Tuning Datasets for Zero-Shot Task Adaptation

### v1

@@ Abstract
We introduce Bonito, an open-source model for conditional task generation: the task of converting unannotated text into task-specific training datasets for instruction tuning. Our goal is to enable zero-shot task adaptation of large language models on users’ specialized, private data. We train Bonito on a new large-scale dataset with 1.65M examples created by remixing existing instruction tuning datasets into meta-templates. The meta-templates for a dataset produce training examples where the input is the unannotated text and the task attribute and the output consists of the instruction and the response. We use Bonito to generate synthetic tasks for seven datasets from specialized domains across three task types—yes-no question answering, extractive question answering, and natural language inference—and adapt language models. We show that Bonito significantly improves the average performance of pretrained and instruction tuned models over the de facto self supervised baseline. For example, adapting Mistral-Instruct-v2 and instruction tuned variants of Mistral and Llama2 with Bonito improves the strong zero-shot performance by 22.1 F1 points whereas the next word prediction objective undoes some of the benefits of instruction tuning and reduces the average performance by 0.8 F1 points. We conduct additional experiments with Bonito to understand the effects of the domain, the size of the training set, and the choice of alternative synthetic task generators. Overall, we show that learning with synthetic instruction tuning datasets is an effective way to adapt language models to new domains. The model, dataset, and code are available at https://github.com/BatsResearch/bonito.
@@ 1 Introduction
Large language models show remarkable zero-shot capabilities by simply learning to predict the next token at scale (Brown et al., 2020; Touvron et al., 2023). By fine-tuning these models on instruction tuning datasets containing many tasks—each comprising an input instruction and a desired response—the model generally improves in its ability to respond to unseen instructions. However, this generalization is still limited by the qualities of the instruction tuning dataset. Existing datasets like the Public Pool of Prompts (P3) (Bach et al., 2022), Natural Instructions (Mishra et al., 2022; Wang et al., 2022), and Dolly-v2 (Conover et al., 2023) are focused on text from the Web, classic natural language datasets, and other tasks that generally do not require specialized domain knowledge, such as biomedical and legal domains. We study how to adapt language models to follow instructions in specialized domains without annotated data.
@@
Self supervision in the form of next word prediction on the target corpus is a simple way to teach language models about new domains (Gururangan et al., 2020). However, this approach requires an enormous amount of training to achieve strong performance (Chen et al., 2023). Further, in our work, we find that self supervision can undo the benefits of instruction tuning (see Section 5.3). Alternatively, continued instruction tuning of models has been shown to improve performance on datasets in specialized domains (Scialom et al., 2022; Shi and Lipani, 2023; Yunxiang et al., 2023; Deng et al., 2023; Singhal et al., 2023a; Wu et al., 2024). However, these works repeat the time-consuming and labor-intensive process of annotating a domain-specific dataset. In this work, we aim to automate the creation of instruction tuning datasets for specialized domains.
@@
We create Bonito, an open-source model to convert unannotated text from specialized domains into task-specific training datasets for instruction tuning (Figure 1). We call this problem conditional task generation. Our key idea is that we can make a new training dataset using existing datasets for instruction tuning. Datasets like P3 (Bach et al., 2022) and the FLAN collection (Longpre et al., 2023) exist as templates that convert semi-structured examples of natural language tasks into a fully prompted format, in which both the input and the desired response are text strings. We start by selecting a subset of the templates in P3 that create tasks from contexts, which are pieces of text that are required for responding to the instruction. For example, a context could be a paragraph that contains a fact or that contains the answer to a question. We also annotate these templates with task attributes, i.e., the type of task they produce. We then use these templates to create meta-templates for training a new language model (see Figure 2). Each meta-template produces training examples in which the input is context and a task attribute, and the output is an entire task: the instruction (including the context) and the desired response. In this way, we can easily create abundant, diverse examples of conditional task generation. We can then train language models on the synthetic datasets to adapt them to the desired task in the target domain.
@@
Bonito significantly improves over self supervision on zero-shot task adaptation of pretrained and instruction tuned models. We use Bonito to generate instruction tuning data for seven datasets across three task types—yes-no question answering (PubMedQA and Privacy Policy QA), extractive question answering (SQuADShifts-NYT, Amazon, and Reddit), and natural language inference (ContractNLI and Vitamin C)—and adapt language models. Our results show that Bonito improved Mistral-7B by 34.7 F1 points and Llama 2 7B by 31.6 F1 points over the self supervised baseline, next word prediction objective. We also consider a more practical setting where we further train Mistral-7B-Instruct-v0.2 and instruction tuned variants of Mistral-7B and Llama 2 7B trained on the T0 split of the P3 dataset. Our results show that Bonito outperforms the strong zero-shot baseline performance by an average of 22.1 F1 points across all the models. On the other hand, we find that self supervision undoes some of the benefits of instruction tuning, i.e., it leads to catastrophic forgetting, resulting in a drop in performance by an average of 0.8 F1 points across all models. Our analysis of Bonito shows that even task specialized models can be further improved by simply learning on Bonito generated tasks (see Section 6.1). We also find that training with more synthetic instructions on datasets like PubMedQA and Vitamin C improves model performance the most compared to other datasets (see Section 6.2). Finally, we perform additional experiments by prompting off-the-shelf open-source models like Zephyr-7B- and Mistral-7B-Instruct-v0.2 and GPT-4 to generate tasks and find they can often improve the pretrained models but still struggle to further increase model performance when they are instruction tuned (see Section 7).
@@ 2 Zero-Shot Task Adaptation
We describe the problem of zero-shot task adaptation. We are given a language model, either pretrained via self supervision or further fine-tuned on a training mixture like P3 (Bach et al., 2022), along with a corpus of unannotated text from the target domain. We also know the target task type e.g., extractive question answering, natural language inference, etc. If the target task type has a fixed set of labels, we assume access to them. Our goal is to adapt the language model to follow task instructions in the target domain without human annotations, i.e., achieve zero-shot task adaptation.
@@ 3 Related Work
Multitask instruction tuning with language models dramatically improves their ability to follow instructions and generalize to new unseen tasks (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022; Longpre et al., 2023; Chung et al., 2022; Zhou et al., 2023; Li et al., 2023). Typically, pretrained models are trained on large-scale training mixtures such as P3 (Bach et al., 2022) and the FLAN collection (Longpre et al., 2023) to follow instructions. In this work, we use P3 to create meta-templates and train Bonito to generate NLP tasks in specialized domains.
@@
Recent work has adapted language models by training on large-scale in-domain datasets(Parmar et al., 2022; Gupta et al., 2022; Singhal et al., 2023b; Deng et al., 2023) or with a few examples from domain-specific tasks (Singhal et al., 2023a). In practice, annotating training datasets for new domains is labor-intensive and expensive. We focus on generating training data for tasks and adapting language models to specialized domains without annotations.
Zero-shot task adaptation is closely related to unsupervised domain adaptation (Ganin and Lempitsky, 2015). In unsupervised domain adaptation, a trained model is used to generate pseudo-labels for the target unlabeled data and then trained on these labels. In our work, naive pseudo-labeling is not applicable as we consider tasks like question answering and natural language inference tasks where a question or a hypothesis is required before predicting the label. Further, popular techniques used in unsupervised domain adaptation such as choosing top-K confident classes (Huang et al., 2022; Menghini et al., 2023) cannot be easily adapted to NLP tasks where there may not be an explicit notion of classes.
There is a growing interest in using retrieval augmented generation (RAG) for open-domain question answering (Lewis et al., 2020; Karpukhin et al., 2020; Siriwardhana et al., 2023). In a RAG pipeline, given a question, the most relevant documents are retrieved before accurately producing an answer with a language model. Our work compliments the RAG pipeline as we assume access to the gold documents or paragraphs from specialized domains and improve the language model’s ability to answer the questions.
@@
Task generation is a fast-growing area of research to adapt large language models to follow instructions (Wang et al., 2023; Taori et al., 2023; Honovich et al., 2023; Köksal et al., 2023). These models condition either GPT or itself on a set of seed task demonstrations and generate new tasks (Wang et al., 2023; Honovich et al., 2023). However, task generation conditioned on the user’s unannotated text has mostly been ignored by these works. Additionally, generating with API-based models is expensive and not usable for proprietary or private research data. On the other hand, Bonito is an open-source model that can be used to create tasks with the user’s unannotated text without additional API costs.
Recently, Li et al. (2023) proposed to learn a backtranslation model, similar to Bonito, to grow and refine their instruction tuning dataset (Gulcehre et al., 2023). However, they focus on generating instructions conditioned on the unannotated text from a web corpus for long-form conversational data where the answer to the instruction is the unannotated text. In contrast, we focus on generating NLP tasks that are conditioned on a task type and unannotated text from a specialized domain. Further, in our experiments, we consider tasks such as question answering and natural language inference that require a question or a hypothesis before generating the appropriate answer.
Concurrent to this work, Yehudai et al. (2024) use in-context learning with Falcon-40B and Llama-65B to generate “grounded tasks” to adapt smaller models like FLAN-T5-XL (3B). These grounded tasks are similar to conditional tasks, except that the instructions do not necessarily refer directly to the user’s text. They might only be based on it, such as asking an open-ended question based on the original text. Our work goes further in several ways. First, we study how to create an open-source model for conditional task generation, as opposed to relying on prompting alone. Second, Bonito has only 7B parameters and we show that it creates data that can improve instruction tuned models of the same size and outperform even larger models like Flan-T5-XXL (11B) (see Appendix D). Third, we evaluate tasks with precise correct/incorrect answers, such as yes-no question answering and natural language inference, as opposed to tasks evaluated with similarity metrics.
@@
Knowledge distillation is a well-studied area (Hinton et al., 2015; Sanh et al., 2019; He et al., 2020). Typically, smaller models learn from the outputs of a larger model. Most recently, API-based models have been used to generate tasks and distilled into smaller models to mimic the abilities of the API-based models (Peng et al., 2023; Gudibande et al., 2023). In this work, we use Bonito to generate tasks based on the user’s context and distill them into pretrained as well as instruction tuned models of the same size for zero-shot task adaptation (see Section 5).
@@
Several works have been proposed in question generation over the years (Mitkov and Ha, 2003; Pan et al., 2020; Lewis et al., 2021; Ushio et al., 2023). They often use heuristics such as templates (Mitkov and Ha, 2003), named entity recognition Lewis et al. (2021), and semantic graphs (Pan et al., 2020). In our work, we train a language model without task-specific heuristics. Ushio et al. (2023) is closely related to our work as they train a unified model to generate extractive questions and answers but only focus on adapting small pretrained language models like T5-Large (770M). In contrast, Bonito can generate tasks beyond extractive question answering and enable zero-shot task adaptation on several task types with large models like Llama 2 7B and Mistral-7B.
@@ 4 Bonito: Learning to Generate Tasks
We describe the steps to create the conditional task generation with attributes dataset to train Bonito. Then, we briefly describe the procedure to create synthetic tasks for the target unannotated texts to adapt language models.
@@
We outline the key properties that we desire in our task generation model: (1) given a corpus containing articles and paragraphs, the model should take the text as input and generate high-quality tasks that require minimal cleaning or post-processing, (2) the model should adhere to the task type like extractive question answering or natural language inference task, and (3) the model should generate diverse tasks for the exact text with varying styles.
Conditional Task Generation with Attributes (CTGA)
To create the model satisfying the key properties, we first create a new training dataset: conditional task generation with attributes (CTGA). The dataset contains 1.65 million examples derived from P3 (Bach et al., 2022) by annotating 323 prompt templates from 39 datasets with 16 task types (see Appendix G).
The prompt templates are used to create the meta-templates, which, in turn, generate the training examples. The meta-template input has a task type (<|tasktype|>) as an attribute followed by the unannotated text or context (<|context|>). The output of the meta-template comprises the attributed task with the prompt or task description and the context ({context}) followed by a pipe symbol (<|pipe|>) and the solution to the task. We use the <|pipe|> symbol to separate the instruction and response pair that is used for adapting the downstream model. Figure 2 shows an input-output example from the CTGA dataset generated using a meta-template.
@@
The dataset is constructed by identifying datasets that require a context to complete the task. For example, SQuAD (Rajpurkar et al., 2016) requires a context to answer extractive question answering tasks whereas CommonSenseQA (Talmor et al., 2019) asks a multiple choice question without providing any relevant text. We identified a total of 39 datasets to be included in CTGA. After selecting relevant datasets with a context, we annotate all the prompts in the dataset with a task type. We annotate a total of 323 prompts with 16 task types. Then, we restructure the prompt template to create the meta-template. Finally, we apply the meta-template to all the examples in a dataset. If the dataset has multiple meta-templates, we uniformly sample one meta-template per example. We limit the total number of examples per dataset to 100,000. The final training dataset is used to train Bonito.
@@
Training the Language Model on the Synthetic Dataset
The trained Bonito model generates synthetic tasks on the target unannotated text for the target task type. For each unannotated text, we generate an instruction and response pair which is then used to train the downstream language model with a cross entropy loss over the output tokens. We provide additional details in Section 5.1.

### v2

@@ Abstract
We introduce Bonito, an open-source model for conditional task generation that converts unannotated text into task-specific training datasets for instruction tuning. We aim to enable zero-shot task adaptation of large language models on users’ specialized, private data. We train Bonito by fine-tuning a pretrained large language model on a new large-scale dataset with 1.65M examples created by remixing existing instruction tuning datasets into meta-templates. The meta-templates for a dataset produce training examples where the input is the unannotated text and the task attribute and the output consists of the instruction and the response. We use Bonito to generate synthetic tasks for seven datasets from specialized domains with unannotated text across three task types—yes-no question answering, extractive question answering, and natural language inference—and adapt language models. We show that Bonito significantly improves the average performance of pretrained and instruction tuned models over the de facto self supervised baseline. For example, adapting Mistral-Instruct-v2 and instruction tuned variants of Mistral and Llama2 with Bonito improves the strong zero-shot performance by 22.1 F1 points whereas the next word prediction objective undoes some of the benefits of instruction tuning and reduces the average performance by 0.8 F1 points. We conduct additional experiments with Bonito to understand the effects of the domain, the size of the training set, and the choice of alternative synthetic task generators. Overall, we show that learning with synthetic instruction tuning datasets is an effective way to adapt language models to new domains. The model, dataset, and code are available at https://github.com/BatsResearch/bonito.
vs
@@ 1 Introduction
Large language models show remarkable zero-shot capabilities by simply learning to predict the next token at scale (Brown et al., 2020; Touvron et al., 2023). By fine-tuning these models on instruction tuning datasets containing many tasks—each comprising an input instruction and a desired response—the model generally improves in its ability to respond to unseen instructions. However, this generalization is still limited by the qualities of the instruction tuning dataset. Existing datasets like Public Pool of Prompts (P3) (Bach et al., 2022), Natural Instructions (Mishra et al., 2022; Wang et al., 2022), and Dolly-v2 (Conover et al., 2023) focus on text from the Web and classic natural language tasks so that they can serve a wide range of use cases, i.e., they are a one-size-fits-all approach. On the other hand, tasks in areas like biomedical and legal domains require specialized, often implicit, domain knowledge. We study how to adapt language models to follow instructions in specialized domains without annotated data.
@@
Self supervision in the form of next word prediction on the target corpus is a simple way to teach language models about new domains (Gururangan et al., 2020). However, this approach requires an enormous amount of training to achieve strong performance (Chen et al., 2023). Further, we find that self supervision can undo the benefits of instruction tuning (see Section 5.3). Alternatively, continued training of models with instructions from specialized domains significantly improves performance (Scialom et al., 2022; Shi and Lipani, 2023; Yunxiang et al., 2023; Deng et al., 2023; Singhal et al., 2023a; Wu et al., 2024). However, they need to repeat the time-consuming and labor-intensive process of annotating a domain-specific dataset. Furthermore, collecting instructions in specialized domains is very expensive because they are annotated by domain experts such as scientists and researchers (Thulke et al., 2024). In this work, we automate the creation of instruction tuning datasets in specialized domains.
@@
We create Bonito, an open-source model to convert unannotated text from specialized domains into task-specific training datasets for instruction tuning (Figure 1). We call this problem conditional task generation. Our key idea is to make a new large-scale dataset called Conditional Task Generation with Attributes (CTGA), to train Bonito, by reorganizing existing instruction tuning datasets (see Figure 2). Instruction tuning datasets like P3 (Bach et al., 2022) exist as templates that convert semi-structured examples of natural language tasks into a fully prompted format, in which both the input and the desired response are text strings. We focus on a subset of the templates in P3 that require a context or a passage to complete the task. For example, a context could be a paragraph that contains a fact or that contains the answer to a question. Then, we remix these templates to create the meta-templates. Each meta-template for a dataset produces training examples in which the input is context and a task attribute such as yes-no question answering, and the output is the entire task: the instruction (including the context) and the desired response. In this way, we can easily create abundant, diverse examples to train Bonito. After training Bonito, we can use new unannotated text from the target domain as the context to generate task-specific synthetic datasets and train specialized language models.
@@
Bonito significantly improves over self supervision on zero-shot task adaptation of pretrained and instruction tuned models. We use Bonito to generate instruction tuning data for seven datasets across three task types—yes-no question answering (PubMedQA and Privacy Policy QA), extractive question answering (SQuADShifts-NYT, Amazon, and Reddit), and natural language inference (ContractNLI and Vitamin C)—and adapt language models. Our results show that Bonito improved Mistral-7B by 34.7 F1 points and Llama 2 7B by 31.6 F1 points over the self supervised baseline, next word prediction objective. We also consider a more practical setting where we further train Mistral-7B-Instruct-v0.2 and instruction tuned variants of Mistral-7B and Llama 2 7B trained on the T0 split of the P3 dataset. Our results show that Bonito outperforms the strong zero-shot baseline performance by an average of 22.1 F1 points across all the models. On the other hand, we find that self supervision undoes some of the benefits of instruction tuning, i.e., it leads to catastrophic forgetting, resulting in a drop in performance by an average of 0.8 F1 points across all models. Our analysis of Bonito shows that even task specialized models can be further improved by simply learning on Bonito generated tasks (see Section 6.1). We also find that training with more synthetic instructions on datasets like PubMedQA and Vitamin C improves model performance the most compared to other datasets (see Section 6.2). We perform additional experiments by prompting off-the-shelf open-source models like Zephyr-7B- and Mistral-7B-Instruct-v0.2 and GPT-4 to generate tasks and find they can often improve the pretrained models but still struggle to increase model performance further when they are instruction tuned (see Section 7). Finally, our human evaluation of Bonito-generated tasks shows that Bonito and GPT-4o generate the same answer on 71 to 77 percent of the tasks. This indicates that Bonito generates high-quality tasks. However, there is room for improvement in generation quality to increase downstream model performance.
@@ 2 Zero-Shot Task Adaptation
We describe the problem of zero-shot task adaptation. We have a language model, either pretrained via self supervision or further fine-tuned on a training mixture like P3 (Bach et al., 2022), along with a corpus of unannotated text from the target domain. We also know the target task type e.g., extractive question answering, natural language inference, etc. If the target task type has a fixed set of labels such as “yes” or “no” in yes-no question answering, we assume access to the label space. Our goal is to adapt the language model to follow task instructions in the target domain without human annotations, achieving zero-shot task adaptation.
@@ 3 Related Work
Multitask instruction tuning of language models dramatically improves their ability to follow instructions and generalize to new unseen tasks (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022; Longpre et al., 2023; Chung et al., 2022; Zhou et al., 2023; Li et al., 2023). Typically, pretrained models are trained to follow instructions on large-scale training mixtures such as P3 (Bach et al., 2022) and the FLAN collection (Longpre et al., 2023). In this work, we use P3 to create meta-templates and train Bonito to generate NLP tasks in specialized domains.
@@
Recent work has adapted language models by training on large-scale in-domain datasets(Parmar et al., 2022; Gupta et al., 2022; Singhal et al., 2023b; Deng et al., 2023) or with a few examples from domain-specific tasks (Singhal et al., 2023a). However, annotating training datasets for new domains is labor-intensive and expensive (Thulke et al., 2024). We focus on generating training datasets in specialized domains and adapting language models without annotations.
Zero-shot task adaptation is closely related to unsupervised domain adaptation (Ganin and Lempitsky, 2015). In unsupervised domain adaptation, a trained model is used to generate pseudo-labels for the target unlabeled data and then trained on these labels. Naive pseudo-labeling cannot be applied to this work since we consider tasks like question answering and natural language inference tasks that require a question or a hypothesis before producing an answer in natural language. Further, popular techniques used in unsupervised domain adaptation such as choosing top-K confident classes (Huang et al., 2022; Menghini et al., 2023) cannot be easily adapted to NLP tasks as there may not be an explicit notion of classes.
There is a growing interest in using retrieval augmented generation (RAG) for domain-specific question answering (Lewis et al., 2020; Karpukhin et al., 2020; Siriwardhana et al., 2023; Zhang et al., 2024). In a RAG pipeline, given a question, the most relevant documents are retrieved before accurately producing an answer with a language model. Our work compliments the RAG pipeline as we assume access to the gold documents or paragraphs from specialized domains and improve the language model’s ability to answer the questions.
@@
Task generation is a fast-growing area of research to adapt large language models to follow instructions (Wang et al., 2023; Taori et al., 2023; Honovich et al., 2023; Köksal et al., 2023; Liu et al., 2024). They typically condition GPT on a set of seed task demonstrations and generate new synthetic tasks (Wang et al., 2023; Honovich et al., 2023). However, task generation conditioned on the user’s unannotated text has mostly been ignored. Additionally, generating with API-based models is expensive and cannot be used for proprietary or private research data (Köksal et al., 2023). On the other hand, Bonito is an open-source model that can be used to create tasks with the user’s unannotated text without additional API costs.
Recently, Li et al. (2023) proposed to learn a backtranslation model, similar to Bonito, to iteratively grow and refine their instruction tuning dataset (Gulcehre et al., 2023). However, they focus on generating instructions conditioned on the unannotated text from a web corpus for long-form conversational data where the answer to the instruction is the unannotated text. In contrast, we focus on generating NLP tasks conditioned on a task type and unannotated text from a specialized domain. Further, we consider tasks such as question answering and natural language inference that require a question or a hypothesis before generating the appropriate answer.
Concurrent to this work, Yehudai et al. (2024) use in-context learning with Falcon-40B and Llama-65B to generate “grounded tasks” to adapt smaller models like FLAN-T5-XL (3B). These grounded tasks are similar to conditional tasks, except the instructions do not necessarily refer directly to the user’s text. They might only be based on it, such as asking an open-ended question based on the original text. Our work goes further in several ways. First, we study how to create an open-source model for conditional task generation, as opposed to relying on prompting alone. Second, Bonito has only 7B parameters and we show that it creates data that can improve instruction tuned models of the same size and outperform even larger models like Flan-T5-XXL (11B) (see Appendix D). Third, we evaluate tasks with precise correct/incorrect answers, such as yes-no question answering and natural language inference, as opposed to tasks evaluated with similarity metrics.
@@
Knowledge distillation is a well-studied area (Hinton et al., 2015; Sanh et al., 2019; He et al., 2020). Typically, smaller models learn from the outputs of a larger model. Most recently, API-based models have been used to generate tasks and distilled into smaller models to mimic the abilities of the API-based models (Peng et al., 2023; Gudibande et al., 2023). In this work, we use Bonito to generate tasks based on the user’s context and distill them into pretrained and instruction tuned models of the same size for zero-shot task adaptation (see Section 5).
@@
Several works have been proposed in question generation over the years (Mitkov and Ha, 2003; Pan et al., 2020; Lewis et al., 2021; Ushio et al., 2023). They often use heuristics such as templates (Mitkov and Ha, 2003), named entity recognition Lewis et al. (2021), and semantic graphs (Pan et al., 2020). In our work, we train a language model without relying on task-specific heuristics. Ushio et al. (2023) is closely related to our work as they train a unified model to generate extractive questions and answers, but only focus on adapting small pretrained language models like T5-Large (770M). In contrast, Bonito can generate tasks beyond extractive question answering and enable zero-shot task adaptation on several task types with large models like Llama 2 7B and Mistral-7B.
@@ 4 Bonito: Learning to Generate Tasks
We describe the steps to create the Conditional Task Generation with Attributes (CTGA) dataset and train Bonito. Then, we briefly describe the procedure to create synthetic instruction tuning datasets with the target unannotated texts to adapt language models.
@@
We outline the key properties that we desire in our conditional task generation model: (1) given a corpus containing articles and paragraphs, the model should take the text as input and generate high-quality tasks that require minimal cleaning or post-processing, (2) the model should adhere to the task type like extractive question answering or natural language inference task, and (3) the model should generate diverse tasks for the exact text with varying styles.
4.1 Creating Bonito: Dataset and Training
To create a model that generates tasks conditioned on text, we create a new training dataset: Conditional Task Generation with Attributes (CTGA). The dataset contains 1.65 million examples derived from P3 (Bach et al., 2022) by annotating 323 prompt templates from 38 datasets with 16 task types (see Appendix H). Then, we train a pretrained large language model on this training dataset to create Bonito.
@@ 4.1 Creating Bonito: Dataset and Training
Figure 2 shows the process of constructing the Conditional Task Generation with Attributes (CTGA) dataset. First, we identify datasets from P3 (Bach et al., 2022) that require a passage or a context to complete the task. For example, SQuAD (Rajpurkar et al., 2016) requires a context to answer extractive question answering tasks, whereas CommonSenseQA (Talmor et al., 2019) asks a multiple choice question without providing any relevant text. We identify a total of 38 datasets to be included in CTGA. For each dataset, we also collect the Jinja111https://jinja.palletsprojects.com/en/3.1.x/ templates from P3. Next, we remix the Jinja templates to create meta-templates. A meta-template is a Jinja template that includes the task attribute or the task type and the key for the context column in the input and the Jinja template for the instruction-response pair in the output with a placeholder {{context}} to avoid repeating the context. Since Jinja templates from P3 do not include a task type, we manually annotate them with a target task type such as yes-no question answering (see Appendix H for details). Overall, we get 323 meta-templates spanning 16 task types (See Table 13 for the list of task types). Finally, we apply the meta-templates to all the examples in a dataset to create the CTGA dataset, i.e., we replace the keys for the columns in the Jinja templates with corresponding key-value pairs from the examples. If the dataset has multiple meta-templates, we uniformly sample one meta-template per example. We limit the total number of examples per dataset to 100,000. The final training dataset is used to train Bonito.
@@
4.2 Adapting Models with Bonito
We use Bonito to create synthetic instruction tuning datasets for the target unannotated texts. Then, the target language model is adapted by training on the synthetic dataset to get the specialized language model.
Generating the Synthetic Dataset
Figure 1 shows the inference with Bonito to generate the synthetic instruction tuning dataset. The unannotated text and the task type are passed to the Bonito model to get the synthetic instruction-response pairs. The process is repeated for all the unannotated text to get the training dataset. The generated pairs are then post-processed into a standardized instruction-response format for instruction tuning. In each generation, we replace {{context}} with the corresponding unannotated text from the input. If the generated output is not parsable due to missing <|pipe|>, we filter them out.
Adapting the Target Model
We train the target language model on the synthetic instruction tuning dataset containing instruction-response pairs. The model is trained using a cross entropy loss over the response tokens. Additional details in Section 5.1.